Sam has 6 rose bushes. He counted the flowers on each of them. There are 8, 2, 5, 4, 11 and 9. Find the Standard Deviation. Is X a “Usual” number of flowers? X is your last two digits of your CUNYFirst ID Number. (00 = 0, 01 = 1, …) (20 points) my id is 03


1
Expert's answer
2021-10-11T13:01:13-0400
"mean=\bar{x}=\dfrac{8+2+5+4+11+9}{6}=6.5"

"Var(x)=\sigma^2=\dfrac{1}{6}((8-6.5)^2+(2-6.5)^2"

"+(5-6.5)^2+(4-6.5)^2+(11-6.5)^2+(9-6.5)^2)"

"=\dfrac{57.5}{6}\approx9.583333"

"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{57.5}{6}}\approx3.0957"

"\mu-\sigma=6.5-\sqrt{\dfrac{57.5}{6}}\approx3.4043"

"\mu+\sigma=6.5+\sqrt{\dfrac{57.5}{6}}\approx9.5957"

By 68–95–99.7 rule

"P(\mu-2\sigma\leq x\leq \mu+2\sigma)=0.9545"

"\mu-2\sigma=6.5-2(\sqrt{\dfrac{57.5}{6}})\approx0.3086"

"\mu+2\sigma=6.5+2(\sqrt{\dfrac{57.5}{6}})\approx12.6914"

"3" is within "[\mu-2\sigma, \mu+2\sigma]." Then  "X=3" is a “Usual” number of flowers.